OpenStudy (anonymous):

Disney World’s Epcot Center has a diameter of 165 feet. Find the approximate surface area and volume of the Epcot Center. Round answers to the nearest whole number and leave pi in the answer. @Directrix

Directrix (directrix):

If the diameter is 165, the the radius is 165/2 = 82.5 A = 4 * π * r² A = 4 * π * (82.5)²

Directrix (directrix):

There are so many people on this site today that the site may crash at any moment. Plus, everything moves slowly, lots of glitches.

Directrix (directrix):

For surface area, did you get 27,225π

OpenStudy (anonymous):

im back

Directrix (directrix):

Look above, did you get that for surface area?

OpenStudy (anonymous):

and yes thats what i got

Directrix (directrix):

V = 4/3 * π * r³ V = 4/3 * π * (82.5)³

OpenStudy (anonymous):

748,688

Directrix (directrix):

I got 748,687.5 π That rounds up to 748,688π

OpenStudy (anonymous):

alright kool

OpenStudy (anonymous):

i got it

Directrix (directrix):

Okay, that's what you got. Look, I wish you'd start adding "pi" to your answers. I hope you don't get on that 5:30 test and leave off the π and miss the answer.

OpenStudy (anonymous):

no it has the pi at the end i just have to click the answer